Which of the following statements is/are correct?

1. A catalyst speeds up a reaction by changing ∆H, the enthalpy of the reaction

2. Addition of a catalyst to a reaction changes the equilibrium constant

3. A catalyst increases the rates of both the forward and reverse reactions

  • .

    3 only

  • B.

    2 only

  • C.

    2 & 3 only

  • D. none

  • E.

    1 & 2 only

Which one of the following statements about the overall order of a reaction is correct?

  • A. the overall order can only be determined by experiment.

  • B.

    the overall order must be a positive integer.

  • C.

    a reaction that is overall second order must be bimolecular.

  • D.

    the overall order can always be determined from the balanced equation for the reaction.

  • E.

    the overall order increases with increasing temperature.

Find the pH of a solution prepared by adding 250 mL of 0.100 M propionic acid to 200 mL of 0.100 M NaOH.

Ka (propionic acid) = 1.3 x 10-5
Answer #1

Correct option: 3 only

Statement 1 is incorrect as a catalyst speeds up a reaction by changing Ea, the activation energy of the reaction.

Statement 2 is incorrect as addition of a catalyst to a reaction has no effect on the equilibrium constant.

Statement 3 is correct as a catalyst increases the rates of both the forward and reverse reactions.

---

Correct option: A. the overall order can only be determined by experiment.

Order of a reaction is determined experimentally.

Option B is incorrect as the overall order of a reaction can be a fraction.

Option C is incorrect as the molecularity and order of a reaction has no correlation.

Option D is incorrect as there is no correlation between order of a reaction and the balanced equation of a reaction.

Option E is incorrect as order of a reaction does not depend on temperature.

---

Moles of propionic acid = 250 mL x 0.100 M = 0.250 mmol

Moles of NaOH = 200 mL x 0.100 M = 0.200 mmol

Now, 0.200 mmol of NaOH will react with 0.250 mmol of propionic acid to give 0.200 mmol of sodium propionate.

Moles of unreacted propionic acid = (0.250 - 0.200) mmol = 0.050 mmol

Total volume of solution after the addition of NaOH = (250 + 200) mL = 450 mL

Hence, the concentration of propionic acid = 0.050 mmol/450 mL = 1.11 x 10-4 M

The concentration of propionate = 0.200 mmol/450 mL = 4.44 x 10-4 M

From Henderson-Hasselbalch equation,

pH = pKa + log([propionate]/[propionic acid])

= - logKa + log(4.44 x 10-4/1.11 x 10-4)

= - log(1.3 x 10-5) + log(4.44 x 10-4/1.11 x 10-4)

= 4.89 + 0.60

= 5.49